# Anisotropy in a wire medium resulting from the rectangularity of a unit cell

### v1

@@ I Introduction
A simple wire medium is formed by parallel metallic wires arranged periodically in a perpendicular plane. Most of the research on this metamaterial [18, 12, 13, 14, 15, 16, 17, 19] has been focused on rectangular/square periodicity, that is, an arrangement of the wires of radii $r_{0}$ in the nodes of a rectangular lattice having the period $a$ in the $x$-direction and the period $b$ in the $y$-direction. In the general case $a$ is not equal to $b$. The geometry of a simple wire medium formed by a rectangular lattice is shown in Fig. 1.
@@ II Anisotropy effect observed
We consider a simple wire medium formed by a rectangular lattice with a $a\times b$ unit cell, where $b$ is assumed to be half as large as $a$. The wires are assumed to be perfectly electrically conducting (PEC) and having the radii $r_{0}=b/20$. We have performed a numerical simulation by applying the periodic conditions corresponding to the wave vector $\vec{q}=(q_{x},q_{y},q_{z})^{T}$ for a unit cell in COMSOL Multyphysics [24]. The isofrequency contours obtained for $q_{z}=0$ are plotted in Fig. 2 with solid lines.
@@
Figure 2 shows that the contours cross the $q_{x}$ and $q_{y}$ axes in the ooints having different coordinates $d_{x}$ and $d_{y}$ ($d_{x}\neq d_{y}$). In other words, the contours arise from the $\Gamma$-point and are elliptical before the contours reach the edges of the Brillouin zone at $q_{x}=\pi/a=\pi/2b$. For the obtained contours the ratio of the ellipse’s semi-axes $d_{x}/d_{y}$ (a contour’s ellipticity) tends to be $\sim 1.13$ in the vicinity of the $\Gamma$-point ($\omega=\omega_{p}+\delta$, $\delta\rightarrow 0$).
The contour ellipticity observed strongly contradicts to the dispersion equation derived in [18, 12]:
@@
that predicts circular contours regardless of the periods ratio $a/b$. The anisotropy for the simple wire medium resulting from the periods of the rectangular lattice changing does not appear to have been ever discussed earlier.
Equation (1) was obtained in [18] by applying the assumption of a small wavenumber within the metamaterial $q$ ($qa,qb\ll\pi$) and a small vacuum wavenumber $k$ ($ka,kb\ll\pi$) in the original (complete) dispersion equation:
@@
We have also verified that a direct numerical solution of the transcendental Eq. (2) matches the results of the computational simulation in COMSOL for the wire media with other ratios $a/b\leq 10$ and other wire radii $b/r_{0}\geq 20$. Based on these results we can suggest that Eq. (2) can be an adequate substitute for the eigenmode solver in COMSOL for a simple wire medium composed of thin wires ($b/r\geq 20$) in the vicinity of the $\Gamma$-point.
@@ III Analytical model of anisotropy
Terms of the first order and cross-terms are equal to zero due to the symmetry of the original Eq. (2). Equation (5) divided by function $F_{0}$ results in a classical form of a quadric surface:
@@
The plasma frequency was calculated by three different means: (i) using COMSOL (the actual plasma frequency), the result is represented by solid lines in the plot; (ii) using Eq. (4), the result is shown in dashed lines; (iii) using Eq. 12, the result is depicted in circular markers. Equation (12) was shown to perform perfectly: the results of numerical modeling match the analytical results. Hereafter the value of $b$ is maintained fixed to make a unified normalization of all the results.
@@ IV Verification of the analytical model
Figure 5: Dependencies of the ellipticity on $a/r_{0}$ ratio for different $a/b$ values. (a) Solid lines plotted by calculating of $q_{x}/q_{y}$, where $q_{x}$ and $q_{y}$ are the solutions of Eq. (2) for fixed $k$ near $k_{p}$. Black dots obtained as $\sqrt{B/A}$, where $A(k,a,b)$ and $B(k,a,b)$ are the functions defined in Eqs. (III) and (III). (b) The same comparision between ellipticities in $yz$-plane obtained via Eq. (2) (solid lines) and by calculating $C(k,a,b)$ and $B(k,a,b)$ functions (III) and (III) (circle markers).
@@
To demonstrate the ability of the low-$q$ model (5) to predict the ellipticity of contours for different configurations of the metamaterial we performed the calculation of $d_{x}/d_{y}$ and $d_{y}/d_{z}$ ratios for a wide range of geometries, see Fig. 5. The circle markers in Fig. 5 were calculated (a) as $\sqrt{B/A}$ using expressions (III-III) and (b) as $\sqrt{C/B}$ by calculating the functions (III-III). The solid lines in both subfigures were obtained by solving Eq. (2) for fixed $k$ three times: (i) for $q_{x}$ with fixed $q_{y}=q_{z}=0$, (ii) for $q_{y}$ with fixed $q_{x}=q_{z}=0$ and (iii) for $q_{z}$ with fixed $q_{x}=q_{y}=0$ at $k$ near $k_{p}$. These two calculations of ellipticity have been shown to perfectly corredpond to each other in both planes: $xy$ and $yz$.
It is important to note, that for extremely thin wires ($b/r\rightarrow\infty$) the shape of isofrequency contours tends to be circular. This is explained by the plasma frequency decreasing with the decrease of wire radii (Fig. 3), which makes the low-$k\&q$ model applicable because of $ka$ and $kb$ becoming small enough. The higher the $a/b$ ratios, the more difficult it is to achieve a small of $ka$ value, hence, the ellipticity for these geometries tends to $1$ for very high $b/r_{0}$ values.

### v2

@@ I Introduction
A simple wire medium is formed by parallel metallic wires arranged periodically in a perpendicular plane. Most of the research on this metamaterial [18, 12, 13, 14, 15, 16, 17, 19] has been focused on rectangular/square periodicity, that is, an arrangement of the wires of radii $r_{0}$ at the nodes of a rectangular lattice having the period $a$ in the $x$-direction and the period $b$ in the $y$-direction. In the general case $a$ is not equal to $b$. The geometry of a simple wire medium formed by a rectangular lattice is shown in Fig. 1.
@@ II Anisotropy effect observed
We consider a simple wire medium formed by a rectangular lattice with an $a\times b$ unit cell, where $b$ is assumed to be half as large as $a$. The wires are assumed to be perfectly electrically conducting (PEC) and having the radii $r_{0}=b/20$. We have performed a numerical simulation by applying the periodic boundary conditions corresponding to the wave vector $\vec{q}=(q_{x},q_{y},q_{z})^{T}$ for a unit cell in COMSOL Multiphysics [24]. The isofrequency contours obtained for $q_{z}=0$ are plotted in Fig. 2 with solid lines.
@@
Figure 2 shows that the contours cross the $q_{x}$ and $q_{y}$ axes in the points having different coordinates $d_{x}$ and $d_{y}$ ($d_{x}\neq d_{y}$). In other words, the contours arise from the $\Gamma$-point and are elliptical before the contours reach the edges of the Brillouin zone at $q_{x}=\pi/a=\pi/2b$. For the obtained contours the ratio of the ellipse’s semi-axes $d_{x}/d_{y}$ (a contour’s ellipticity) tends to be $\sim 1.13$ in the vicinity of the $\Gamma$-point ($\omega=\omega_{p}+\delta$, $\delta\rightarrow 0$).
The electromagnetic anisotropy for the simple wire medium resulting from the periods of the rectangular lattice changing does not appear to have been ever discussed earlier. However, an analogous [25] acoustic task for a two-dimensional array of rigid cylinders was reported in [26], where an anisotropy for an acoustic wave was analytically and numerically shown in the long wavelength limit.
The observed contour ellipticity can not be described by the dispersion equation derived in [18, 12]:
@@
that predicts circular contours regardless of the periods ratio $a/b$. Equation (1) was obtained in [18] by applying the assumption of a small wavenumber within the metamaterial $q$ ($qa,qb\ll\pi$) and a small vacuum wavenumber $k$ ($ka,kb\ll\pi$) in the original (complete) dispersion equation:
@@
We have also verified that a direct numerical solution of the transcendental Eq. (2) matches the results of the computational simulation in COMSOL for the wire media with other ratios $a/b\leq 10$ and other wire radii $b/r_{0}\geq 20$. Based on these results we can suggest that Eq. (2) can be an adequate substitute for the eigenmode solver in COMSOL for a simple wire medium composed of thin wires ($b/r_{0}\geq 20$) in the vicinity of the $\Gamma$-point.
@@ III Analytical model of anisotropy
Terms of the first order and cross-terms are equal to zero due to the symmetry of the metamaterial geometry. Equation (5) divided by function $F_{0}$ results in a classical form of a quadric surface:
@@
The plasma frequency was calculated by three different means: (i) using COMSOL (the actual plasma frequency), the result is represented by solid lines in the plot; (ii) using Eq. (4), the result is shown in dashed lines; (iii) using Eq. (12), the result is depicted in circular markers. Equation (12) was shown to perform perfectly: the results of numerical modeling match the analytical results. Hereafter the value of $b$ is maintained fixed to make a unified normalization of all the results.
@@ IV Verification of the analytical model
Figure 5: Dependencies of the ellipticity on $b/r_{0}$ ratio for different $a/b$ values. (a) Solid lines plotted by calculating of $q_{x}/q_{y}$, where $q_{x}$ and $q_{y}$ are the solutions of Eq. (2) for fixed $k$ near $k_{p}$. Black dots obtained as $\sqrt{B/A}$, where $A(k,a,b)$ and $B(k,a,b)$ are the functions defined in Eqs. (III) and (III). (b) The same comparision between ellipticities in $yz$-plane obtained via Eq. (2) (solid lines) and by calculating $C(k,a,b)$ and $B(k,a,b)$ functions (III) and (III) (circle markers).
@@
To demonstrate the ability of the low-$q$ model (5) to predict the ellipticity of contours for different configurations of the metamaterial we performed the calculation of $d_{x}/d_{y}$ and $d_{y}/d_{z}$ ratios for a wide range of geometries, see Fig. 5. The circle markers in Fig. 5 were calculated (a) as $\sqrt{B/A}$ using expressions (III-III) and (b) as $\sqrt{C/B}$ by calculating the functions (III-III). The solid lines in both subfigures were obtained by solving Eq. (2) for fixed $k$ three times: (i) for $q_{x}$ with fixed $q_{y}=q_{z}=0$, (ii) for $q_{y}$ with fixed $q_{x}=q_{z}=0$ and (iii) for $q_{z}$ with fixed $q_{x}=q_{y}=0$ at $k$ near $k_{p}$. These two calculations of ellipticity have been shown to perfectly correspond to each other in both planes: $xy$ and $yz$.
It is important to note, that for extremely thin wires ($b/r_{0}\rightarrow\infty$) the shape of isofrequency contours tends to be circular. This is explained by the plasma frequency decreasing with the decrease of wire radii (Fig. 3), which makes the low-$k\&q$ model applicable because of $ka$ and $kb$ becoming small enough. The higher the $a/b$ ratios, the more difficult it is to achieve a small $ka$ value, hence, the ellipticity for these geometries tends to $1$ for very high $b/r_{0}$ values.